### v1

@@ 8. Quantitative Khintchine dichotomy in $\mathbb{R}^{d}$ from equidistribution
We show that an arbitrary probability measure $\xi$ on $\mathbb{R}^{d}$ obeys the Khintchine dichotomy provided that the pushfoward $a(t)u(\boldsymbol{s})\operatorname{SL}_{d+1}(\mathbb{Z})\,\mathrm{d}\xi(\boldsymbol{s})$ satisfies certain effective equidistribution properties in $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ for large $t$. Combined with Proposition˜7.3, this yields ˜1.1.
@@
(8.1) $\forall f\in B^{\infty}_{\infty,l}(X),\,\forall t>1,\\ \left\lvert\int_{\mathbb{R}^{d}}f\bigl{(}a(t)u(\boldsymbol{s})x_{0}\bigr{)}\,\mathrm{d}\xi(\boldsymbol{s})-m_{X}(f)\right\rvert\leq C{\mathcal{S}}_{\infty,l}(f)t^{-c}.$
@@
(8.2) $\forall f_{1},f_{2}\in B^{\infty}_{\infty,l}(X),\,\forall t_{1}>1,\,\forall t_{2}>t_{1}^{1+\eta},\\ \Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}.$
@@
(8.3) $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{2,l}(f_{1}){\mathcal{S}}_{2,l}(f_{2})t_{1}^{c}t_{2}^{-c}\\ +C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}.$
The implication from (8.2) to (8.3) (again, parameters $C,c,l$ may differ) is explained in [7, Section 7.1]. The idea is that (8.2) implies single equidistribution, which in turn, by decay of matrix coefficients, yields effective double equidistribution in the short range regime $t_{1}\leq t_{2}\leq t_{2}^{1+\eta}$ for sufficiently small $\eta$.
@@
Theorem 8.2 (Convergent case [28, Theorem 9.1]).
@@
(8.4) $\big{|}\{(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\llbracket 1,N\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\}\big{|}\sim_{\boldsymbol{s},\psi}\sum_{q=1}^{N}\psi(q)^{d}.$
@@
(8.5) $\big{|}\{(\boldsymbol{p},q)\in{\mathcal{P}}(\mathbb{Z}^{d+1})_{N}\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\}\big{|}\sim_{\boldsymbol{s},\psi}\zeta(d+1)^{-1}\sum_{q=1}^{N}\psi(q)^{d}.$
@@
Note also that in both cases (non-primitive and primitive), given an arbitrary subset of subscripts $I\subseteq\llbracket 1,d\rrbracket$, we may replace the condition $0\leq qs_{i}-p_{i}<\psi(q)$ ($i\in I$) in the above sets by $0\leq p_{i}-qs_{i}<\psi(q)$ ($i\in I$) without affecting the asymptotic.
The work conducted in previous sections guarantees that self-similar measures satisfy the conditions required in Theorems 8.2, 8.3. Provided ˜8.3 holds, we directly deduce ˜1.1.
@@
The convergent case follows from combining Theorems 1.2, 8.2. The divergent case is a consequence of Proposition˜7.3 and ˜8.3 (along with its subsequent remark). ∎
It remains to show ˜8.3. We will distinguish the cases where $d\geq 2$ and $d=1$. The reason for that distinction is that the Siegel transform of the characteristic function of a ball in $\mathbb{R}^{d}$ has finite second moment when $d\geq 2$ and infinite second moment when $d=1$. We present the case $d\geq 2$ and explain afterward how the proof can be adapted, by the mean of a suitable truncation, to obtain the case $d=1$.
@@ 8.1. Case $d\geq 2$, the lower bound
We show the lower asymptotic in ˜8.3.
@@
$\mathscr{T}_{N}(\boldsymbol{s}):=\big{|}\{(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\llbracket 1,N\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\}\big{|}.$
@@
$\mathscr{S}_{k}(\boldsymbol{s}):=\big{|}\{(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\rrbracket\tau^{k-1},\tau^{k}\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(\tau^{k})\}\big{|}.$
@@
(8.6) $\mathscr{T}_{N}(\boldsymbol{s})\geq\mathscr{T}_{\tau^{n}}(\boldsymbol{s})\geq\sum_{k=1}^{n}\mathscr{S}_{k}(\boldsymbol{s}).$
@@
Under the assumptions of ˜8.3 and with $d\geq 2$, we have for every $\tau\in(1,2]$, for $\xi$-almost all $\boldsymbol{s}\in\mathbb{R}^{d}$, for every $\eta>0$, for all large enough $n$,
$\sum_{k=1}^{n}\mathscr{S}_{k}(\boldsymbol{s})\geq(1-\tau^{-1}-\eta)\sum_{k=1}^{n}\psi(\tau^{k})^{d}\tau^{k}.$
The lower bound in ˜8.3 follows directly:
Proof of lower bound in (8.4) using Proposition˜8.4.
@@
$(1-\tau^{-1})\tau\sum_{k=1}^{n}\psi(\tau^{k})^{d}\tau^{k}\geq(1-2\varepsilon)\sum_{q=1}^{N}\psi(q)^{d}$
Choose $\tau$ close enough to $1$ so that $\tau^{-1}\geq(1-\varepsilon)$. Choose $\eta>0$ with $\eta\lll_{\tau,\varepsilon}1$ so that $(1-\tau^{-1}-\eta)\geq(1-\varepsilon)(1-\tau^{-1})$. Using Proposition˜8.4, then (8.6), we obtain that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $N$,
@@
We now focus on the proof of Proposition˜8.4. For that, we need a dynamical interpretation of $\mathscr{S}_{k}(\boldsymbol{s})$. Recall $X=\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ throughout the section. Given a measurable non-negative function $f:\mathbb{R}^{d+1}\rightarrow\mathbb{R}_{\geq 0}$, we denote by $\widetilde{f}:X\rightarrow[0,+\infty]$ its Siegel transform. It is given by: for $g\in G$,
@@
$\psi(\tau^{k})=r_{k}t_{k}^{-\frac{1}{d+1}},\quad\tau^{k}=t_{k}^{\frac{d}{d+1}}r_{k},$
@@
(8.8) $\mathscr{S}_{k}(\boldsymbol{s})=\widetilde{\mathbbm{1}}_{R_{k}}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
Let $\gamma_{1}\in(0,1/2)$ be a small parameter to be specified later. We partition the subscripts $k$’s into two families : $\operatorname{K_{big}}:=\{k\geq 3\,:\,r_{k}>t_{k}^{\gamma_{1}}\}$, and $\operatorname{K_{small}}:=\mathbb{N}_{\geq 3}\smallsetminus\operatorname{K_{big}}$. Given $n\geq 1$, we set $\operatorname{K_{big}}(n)=\operatorname{K_{big}}\cap\llbracket 1,n\rrbracket$, and $\operatorname{K_{small}}(n)=\operatorname{K_{small}}\cap\llbracket 1,n\rrbracket$. We will establish the lower asymptotics required by Proposition˜8.4 for the sums $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ and $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ separately.
@@
For $r>1$, let $f_{r}:X\rightarrow[0,1]$ be a smooth function such that ${\mathcal{S}}_{\infty,l}(f_{r})\ll 1$ and ${\mathbbm{1}}_{\operatorname{dist}(\cdot,x_{0})\geq r}\leq f_{r}\leq{\mathbbm{1}}_{\operatorname{dist}(\cdot,x_{0})\geq r/2}$. Then by Lemma˜4.5,
@@
$\xi\{\boldsymbol{s}\,:\,\operatorname{dist}(a(t)u(\boldsymbol{s})x_{0},x_{0})\geq r\}\leq\int_{\mathbb{R}^{d}}f_{r}(a(r)u(\boldsymbol{s})x_{0})\,\mathrm{d}\xi(\boldsymbol{s})\ll Me^{-r/M}+t^{-c}.$
@@
The next lemma expresses that the counting measure on a covolume $1$ lattice of $\mathbb{R}^{d+1}$ is a good volume estimate for a box in $\mathbb{R}^{d+1}$, provided the box has large enough sidelength depending on the distorsion of the lattice.
@@
We infer from Lemmas 8.5, 8.6 the asymptotic lower bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$.
@@
Recall that for every $\boldsymbol{s}\in\mathbb{R}^{d}$, $k\geq 3$, we have $\mathscr{S}_{k}(\boldsymbol{s})=\widetilde{{\mathbbm{1}}}_{R_{k}}(a(t_{k})u(\boldsymbol{s})x_{0})$. Assuming $k\in\operatorname{K_{big}}$, we have that $R_{k}$ is a box of minimal sidelength $(1-\tau^{-1})r_{k}\geq(1-\tau^{-1})t^{\gamma_{1}}_{k}$. Moreover Lemma˜8.5 guarantees that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $k$, say $k\geq k_{\boldsymbol{s}}$, we have $a(t_{k})u(\boldsymbol{s})\in g\operatorname{SL}_{d+1}(\mathbb{Z})$ where $\|g\|\leq(\log t_{k})^{O(1)}\leq t_{k}^{\gamma_{1}/100}$. Invoking Lemma˜8.6, we obtain in this context
@@
(8.9) $\chi_{k}(x)=\begin{cases}1,\quad\text{if }\operatorname{inj}(x)\geq t_{k}^{-\gamma_{2}},\\ 0,\quad\text{ otherwise.}\end{cases}$
@@
We first recall well-known moment estimates for the Siegel transform of characteristic functions of bounded sets, see [50, pages 2-3]. We emphasize here that we work under the assumption $d\geq 2$ (otherwise (8.11) does not hold, see §8.3).
Fact 8.8 (Moments of Siegel transforms [50]).
@@
(8.10) $\int_{X}\widetilde{{\mathbbm{1}}}_{R}\,\mathrm{d}m_{X}=\operatorname{Leb}(R)$
(8.11) $\int_{X}(\widetilde{{\mathbbm{1}}}_{R})^{2}\,\mathrm{d}m_{X}=\operatorname{Leb}(R)^{2}+O(\operatorname{Leb}(R)).$
@@
We deduce from ˜8.8 and Lemma˜8.9 several moment estimates for the functions $\varphi_{k}$.
@@
(8.12) $m_{X}(\varphi_{k})=\operatorname{Leb}(R_{k}^{-})-O(t_{k}^{-\gamma_{2}/M}),$
(8.13) ${\mathcal{S}}_{\infty,l}(\varphi_{k})\ll t_{k}^{M\gamma_{2}},$
(8.14) ${\mathcal{S}}_{2,l}(\varphi_{k})\ll m_{X}(\varphi_{k})+\sqrt{m_{X}(\varphi_{k})}+t_{k}^{-\gamma_{2}/M}.\$
@@
Let us prove (8.12). Note first $m_{X}(\varphi_{k})=m_{X}(\chi_{k}\widetilde{{\mathbbm{1}}}_{R_{k}^{-}})$. Applying (8.10), followed by the Cauchy-Schwarz inequality, Lemma˜4.5 and (8.11), we find
@@
$=\sqrt{m_{X}\{\operatorname{inj}<t^{-\gamma_{2}}_{k}\}}\,\|\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}\|_{L^{2}}$
@@
Finally, we check (8.14). By Lemma˜8.9 followed by (8.11), we have
@@
$Y_{k}:\mathbb{R}^{d}\to\mathbb{R},\quad\boldsymbol{s}\mapsto\varphi_{k}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
@@
(8.15) $=\int_{X}(\varphi_{k}(x)-y_{k})^{2}\,\mathrm{d}m_{X}(x)+O({\mathcal{S}}_{\infty,l}([\varphi_{k}-y_{k}]^{2})t_{k}^{-c}).$
@@
We now estimate the main term of (8.15). By expanding the square, then using Lemma˜8.9 and $m_{G}(\theta)=1$, we see that
@@
Using (8.11), (8.12) and $y_{k}\geq 1$, the main term is bounded by $O(y_{k})$. The result follows. ∎
@@
(8.16) $\psi(q)\geq q^{-1/d}\log^{-2/d}(q),\quad\forall q\in\operatorname{K_{small}}.$
@@
$\mathbb{E}[(\sum_{j\in J}Z_{j})^{2}]\ll\left(1+\sum_{j\in J}y_{j}\right)^{3/2}.$
@@
$Y_{J}:=\sum_{k\in J}Y_{k},\quad\quad\quad\quad y_{J}:=\sum_{k\in J}y_{k},\quad\quad\quad\quad Z_{J}=Y_{J}-y_{J}.$
@@
(8.17) $\mathbb{E}[Z_{j}Z_{k}]=\mathbb{E}[Y_{j}Y_{k}]-y_{j}y_{k}-\mathbb{E}[Z_{j}]y_{k}-y_{j}\mathbb{E}[Z_{k}].$
@@
(8.18) $\ll\sqrt{y_{j}y_{k}}t_{j}^{c}t_{k}^{-c}+y_{k}t_{j}^{-c/2}+t_{k}^{-c/2}+t_{j}^{c-\gamma_{2}/M}t_{k}^{-c}$
where the second inequality assumes $\gamma_{2}\leq c/(4M)$.
@@
(8.19) $\lvert\mathbb{E}[Z_{j}]\rvert\ll t_{j}^{-c+M\gamma_{2}}\ll t_{j}^{-c/2}.$

### v2

@@ 9. Quantitative Khintchine dichotomy in $\mathbb{R}^{d}$ from equidistribution
We show that an arbitrary probability measure $\xi$ on $\mathbb{R}^{d}$ obeys the Khintchine dichotomy provided that the pushfoward $a(t)u(\boldsymbol{s})\operatorname{SL}_{d+1}(\mathbb{Z})\,\mathrm{d}\xi(\boldsymbol{s})$ satisfies certain effective equidistribution properties in $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ for large $t$. Combined with Proposition˜8.3, this yields ˜1.1.
@@
$\forall f\in B^{\infty}_{\infty,l}(X),\,\forall t>1,\\ \left\lvert\int_{\mathbb{R}^{d}}f\bigl(a(t)u(\boldsymbol{s})x_{0}\bigr)\,\mathrm{d}\xi(\boldsymbol{s})-m_{X}(f)\right\rvert\leq C{\mathcal{S}}_{\infty,l}(f)t^{-c}.$ (9.1)
@@
$\forall f_{1},f_{2}\in B^{\infty}_{\infty,l}(X),\,\forall t_{1}>1,\,\forall t_{2}>t_{1}^{1+\eta},\\ \Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}$ (9.2)
@@
$\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{2,l}(f_{1}){\mathcal{S}}_{2,l}(f_{2})t_{1}^{c}t_{2}^{-c}\\ +C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}.$ (9.3)
The implication from (9.2) to (9.3) (again, parameters $C,c,l$ may differ) is explained in [3, Section 7.1]. The idea is that (9.2) implies single equidistribution, which in turn, by decay of matrix coefficients, yields effective double equidistribution in the short range regime $t_{1}\leq t_{2}\leq t_{2}^{1+\eta}$ for sufficiently small $\eta$.
@@
Theorem 9.2 (Convergent case [36, Theorem 9.1]).
@@
$\big|\{\,(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\llbracket 1,N\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\,\}\big|\sim_{\boldsymbol{s},\psi}\sum_{q=1}^{N}\psi(q)^{d}.$ (9.4)
@@
$\big|\{(\boldsymbol{p},q)\in{\mathcal{P}}(\mathbb{Z}^{d+1})_{N}\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\}\big|\sim_{\boldsymbol{s},\psi}\zeta(d+1)^{-1}\sum_{q=1}^{N}\psi(q)^{d},$ (9.5)
@@
Note also that in both cases (non-primitive and primitive), given an arbitrary subset of subscripts $I\subseteq\llbracket 1,d\rrbracket$, we may replace the conditions $0\leq qs_{i}-p_{i}<\psi(q)$ ($i\in I$) in the above sets by $0\leq p_{i}-qs_{i}<\psi(q)$ ($i\in I$) without affecting the asymptotic.
The work conducted in previous sections guarantees that self-similar measures satisfy the conditions required in Theorems 9.2, 9.3. Provided ˜9.3 holds, we directly deduce ˜1.1.
@@
The convergent case follows from combining Theorems 1.2, 9.2. The divergent case is a consequence of Proposition˜8.3 and ˜9.3 (along with its subsequent remark). ∎
It remains to show ˜9.3. We will distinguish the cases where $d\geq 2$ and $d=1$. The reason for that distinction is that the Siegel transform of the characteristic function of a ball in $\mathbb{R}^{d}$ has finite second moment when $d\geq 2$ and infinite second moment when $d=1$. We present the case $d\geq 2$ and explain afterward how the proof can be adapted, by the mean of a suitable truncation, to obtain the case $d=1$.
@@ 9.1. Case $d\geq 2$, the lower bound
We show the lower asymptotic in ˜9.3.
@@
$\mathscr{T}_{N}(\boldsymbol{s}):=\bigl\lvert\{\,(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\llbracket 1,N\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\,\}\bigr\rvert$.
@@
$\mathscr{S}_{k}(\boldsymbol{s}):=\bigl\lvert\{\,(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\rrbracket\tau^{k-1},\tau^{k}\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(\tau^{k})\,\}\bigr\rvert$.
@@
$\mathscr{T}_{N}(\boldsymbol{s})\geq\mathscr{T}_{\tau^{n}}(\boldsymbol{s})\geq\sum_{k=1}^{n}\mathscr{S}_{k}(\boldsymbol{s}).$ (9.6)
@@
Under the assumptions of ˜9.3 and with $d\geq 2$, we have for every $\tau\in(1,2]$, for $\xi$-almost all $\boldsymbol{s}\in\mathbb{R}^{d}$, for every $\eta>0$, for all large enough $n$,
$\sum_{k=1}^{n}\mathscr{S}_{k}(\boldsymbol{s})\geq(1-\tau^{-1}-\eta)\sum_{k=1}^{n}\psi(\tau^{k})^{d}\tau^{k}.$ (9.7)
The lower bound in ˜9.3 follows directly:
Proof of the lower bound in (9.4) using Proposition˜9.4.
@@
$(1-\tau^{-1})\tau\sum_{k=1}^{n}\psi(\tau^{k})^{d}\tau^{k}\geq(1-2\varepsilon)\sum_{q=1}^{N}\psi(q)^{d}$ (9.8)
We now compare the left-hand side of (9.8) with the right-hand side of (9.7). For that, choose $\tau$ close enough to $1$ so that $\tau^{-1}\geq 1-\varepsilon$. Taking $\eta>0$ with $\eta\lll_{\tau,\varepsilon}1$, we then have $(1-\tau^{-1}-\eta)\geq(1-\varepsilon)^{2}(1-\tau^{-1})\tau$. Using Proposition˜9.4, then (9.6), we obtain that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $N$,
@@
We now focus on the proof of Proposition˜9.4. For that, we need a dynamical interpretation of $\mathscr{S}_{k}(\boldsymbol{s})$. Recall $X=\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ throughout the section. Given a measurable non-negative function $f:\mathbb{R}^{d+1}\rightarrow\mathbb{R}_{\geq 0}$, we denote by $\widetilde{f}:X\rightarrow[0,+\infty]$ its Siegel transform. It is given by: for $g\in G$,
@@
$\psi(\tau^{k})=r_{k}t_{k}^{-\frac{1}{d+1}},\qquad\tau^{k}=t_{k}^{\frac{d}{d+1}}r_{k},$
@@
$\mathscr{S}_{k}(\boldsymbol{s})=\widetilde{\mathbbm{1}}_{R_{k}}\bigl(a(t_{k})u(\boldsymbol{s})x_{0}\bigr).$ (9.10)
Let $\gamma_{1}\in(0,1/2)$ be a small parameter to be specified later. We partition the subscripts $k$’s into two families :
$\operatorname{K_{big}}:=\{\,k\geq 3\,:\,r_{k}>t_{k}^{\gamma_{1}}\,\}\qquad\text{and}\qquad\operatorname{K_{small}}:=\{\,k\geq 3\,:\,k\notin\operatorname{K_{big}}\,\}.$
Given $n\geq 1$, we set $\operatorname{K_{big}}(n)=\operatorname{K_{big}}\cap\llbracket 1,n\rrbracket$, and $\operatorname{K_{small}}(n)=\operatorname{K_{small}}\cap\llbracket 1,n\rrbracket$. We will establish the lower asymptotics required by Proposition˜9.4 for the sums $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ and $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ separately.
@@
For $r>1$, let $f_{r}:X\rightarrow[0,1]$ be a smooth function such that ${\mathcal{S}}_{\infty,l}(f_{r})\ll 1$ and ${\mathbbm{1}}_{\operatorname{dist}(\cdot,x_{0})\geq r}\leq f_{r}\leq{\mathbbm{1}}_{\operatorname{dist}(\cdot,x_{0})\geq r/2}$. Then by Lemma˜5.5,
@@
$\xi\{\boldsymbol{s}\,:\,\operatorname{dist}(a(t)u(\boldsymbol{s})x_{0},x_{0})\geq r\}\leq\int_{\mathbb{R}^{d}}f_{r}(a(t)u(\boldsymbol{s})x_{0})\,\mathrm{d}\xi(\boldsymbol{s})\ll Me^{-r/M}+t^{-c}.$
@@
The next lemma expresses that the counting measure on a covolume $1$ lattice of $\mathbb{R}^{d+1}$ is a good volume estimate for a box in $\mathbb{R}^{d+1}$, provided the box has large enough side length depending on the distortion of the lattice.
@@
We infer from Lemmas 9.5, 9.6 the asymptotic lower bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$.
@@
Recall that for every $\boldsymbol{s}\in\mathbb{R}^{d}$, $k\geq 3$, we have $\mathscr{S}_{k}(\boldsymbol{s})=\widetilde{{\mathbbm{1}}}_{R_{k}}(a(t_{k})u(\boldsymbol{s})x_{0})$. Assuming $k\in\operatorname{K_{big}}$, we have that $R_{k}$ is a box of minimal side length $(1-\tau^{-1})r_{k}\geq(1-\tau^{-1})t^{\gamma_{1}}_{k}$. Moreover Lemma˜9.5 guarantees that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $k$, say $k\geq k_{\boldsymbol{s}}$, we have $a(t_{k})u(\boldsymbol{s})\in g\operatorname{SL}_{d+1}(\mathbb{Z})$ where $\|g\|\leq(\log t_{k})^{O(1)}\leq t_{k}^{\gamma_{1}/100}$. Invoking Lemma˜9.6, we obtain in this context
@@
$\chi_{k}(x)=\begin{cases}1\quad\text{if }\operatorname{inj}(x)\geq t_{k}^{-\gamma_{2}},\\ 0\quad\text{ otherwise.}\end{cases}$ (9.11)
@@
We first recall well-known moment estimates for the Siegel transform of characteristic functions of bounded sets, see [58, pages 2-3]. We emphasize here that we work under the assumption $d\geq 2$ (otherwise (9.13) does not hold, see §9.3).
Fact 9.8 (Moments of Siegel transforms [58]).
@@
$\int_{X}\widetilde{{\mathbbm{1}}}_{R}\,\mathrm{d}m_{X}=\operatorname{Leb}(R),$ (9.12)
$\int_{X}(\widetilde{{\mathbbm{1}}}_{R})^{2}\,\mathrm{d}m_{X}=\operatorname{Leb}(R)^{2}+O(\operatorname{Leb}(R)).$ (9.13)
@@
We deduce from ˜9.8 and Lemma˜9.9 several moment estimates for the functions $\varphi_{k}$.
@@
$m_{X}(\varphi_{k})=\operatorname{Leb}(R_{k}^{-})-O(t_{k}^{-\gamma_{2}/M}),$ (9.14)
${\mathcal{S}}_{\infty,l}(\varphi_{k})\ll t_{k}^{M\gamma_{2}},$ (9.15)
${\mathcal{S}}_{2,l}(\varphi_{k})\ll m_{X}(\varphi_{k})+\sqrt{m_{X}(\varphi_{k})}+t_{k}^{-\gamma_{2}/M}.\$(9.16)
@@
Let us prove (9.14). Note first $m_{X}(\varphi_{k})=m_{X}(\chi_{k}\widetilde{{\mathbbm{1}}}_{R_{k}^{-}})$. Applying (9.12), followed by the Cauchy-Schwarz inequality, Lemma˜5.5 and (9.13), we find
@@
$\leq\sqrt{m_{X}\{\operatorname{inj}<t^{-\gamma_{2}}_{k}\}}\,\|\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}\|_{L^{2}}$
@@
Finally, we check (9.16). By Lemma˜9.9 followed by (9.13), we have
@@
$Y_{k}:\mathbb{R}^{d}\to\mathbb{R},\qquad\boldsymbol{s}\mapsto\varphi_{k}\bigl(a(t_{k})u(\boldsymbol{s})x_{0}\bigr).$
@@
$=\int_{X}(\varphi_{k}(x)-y_{k})^{2}\,\mathrm{d}m_{X}(x)+O({\mathcal{S}}_{\infty,l}([\varphi_{k}-y_{k}]^{2})t_{k}^{-c}).$ (9.17)
@@
We now estimate the main term of (9.17). By expanding the square, then using Lemma˜9.9 and $m_{G}(\theta)=1$, we see that
@@
Using (9.13), (9.14) and $y_{k}\geq 1$, we obtain that the main term is bounded by $O(y_{k})$. The result follows. ∎
@@
$\psi(q)\geq q^{-1/d}\log^{-2/d}(q),\qquad\forall q\in\operatorname{K_{small}}.$ (9.18)
@@
$\mathbb{E}\left[(\sum_{j\in J}Z_{j})^{2}\right]\ll\left(1+\sum_{j\in J}y_{j}\right)^{3/2}.$
@@
$Y_{J}:=\sum_{k\in J}Y_{k},\qquad\qquad\qquad y_{J}:=\sum_{k\in J}y_{k}\qquad\qquad\text{and}\qquad\qquad Z_{J}=Y_{J}-y_{J}.$
@@
$\mathbb{E}[Z_{j}Z_{k}]=\mathbb{E}[Y_{j}Y_{k}]-y_{j}y_{k}-\mathbb{E}[Z_{j}]y_{k}-y_{j}\mathbb{E}[Z_{k}].$ (9.19)
@@
$\ll\sqrt{y_{j}y_{k}}t_{j}^{c}t_{k}^{-c}+y_{k}t_{j}^{-c/2}+t_{k}^{-c/2}+t_{j}^{c-\gamma_{2}/M}t_{k}^{-c}$ (9.20)
where the second inequality assumes $\gamma_{2}\leq c/(4M)$, and uses that $t_{j}\leq t_{k}$.
@@
$\lvert\mathbb{E}[Z_{j}]\rvert\ll t_{j}^{-c+M\gamma_{2}}\ll t_{j}^{-c/2}.$ (9.21)